The equilibrium concentrations of in a gas mixture at a particular temperature are 0.13 mol. dm-3 H2, 0.70 mol. dm-3 I2, and 2.1 mol. dm-3 HI. What equilibrium concentrations are obtained when 0.200 mol of HI is injected into an empty 500 cm3 container?

1
Expert's answer
2022-01-31T20:04:02-0500

"H_2+I_2\rightleftharpoons2HI"


"K_c=\frac{[HI]^2}{[H_2][I_2]}"


"K_c=\frac{(2.1)^2}{0.13×0.7}=48.46"


"[HI]=\frac{0.2}{0.5}=0.4m"


"\begin{matrix}
 & &H_2&&I_2&&2HI \\
 I&&0&&0&&0.4 \\
C&&x&&x&&-2x\\
E&&x&&x&&0.4-2x
\end{matrix}"



"48.46=\frac{(0.4-2x)^2}{(x)(x)}"


"6.961=\frac{0.4-2x}{x}"


"x=0.\n04464"


"[HI]=0.4-2(0.04464)"

"=0.31072"


"[H_2]=[I_2]=0.04464"